Circular No. 6244

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995ag IN UGC 11861 J. Mueller reports her discovery on a supernova mag ∼ 17 located

20′′ west and 25′′ south of the center of UGC 11861 (α = 21^{h}56^{m}.4, δ = +73^{o}15′, equinox 2000.0); it was found on a IV-N plate taken by D. Griffith and Mueller on Sept. 28 UT with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. No object is present at the position of SN 1995ag on original Sky Survey prints or on an F plate taken by C. Brewer and J. D. Mendenhall on 1992 June 28. A spectrum obtained on Oct. 16 at the 5-m Hale telescope (+ double spectrograph) by I. N. Reid and J. Gizis confirms this to be a type-I supernova.

COMET C/1995 O1 (HALE-BOPP) M. F. A’Hearn, M. Haken, and P. D. Feldman report observations

of this comet with the International Ultraviolet Explorer satellite in late August and early September: “The upper limit on water vaporization is 10^{29} s^{−1}, which is not low enough to provide an interesting constraint on the radius of the nucleus. The dust production corresponds to a value of Afρ = 400.00 m, one of the largest values ever observed (A is the albedo of the grains, f is the filling factor of the grains, and ρ is the radius of the field of view; cf. A’Hearn et al. 1984, A.J. 89, 579). Further details are available on the PDS-SBN Hale-Bopp Bulletin Board at URL

http://pdssbn.astro.umd.edu.”

Regarding his item on IAUC 6240, M. Kidger adds: “The new jet was found by R. Torres-Chico. Further analysis of those images, as well as more images taken on Oct. 1 and 2 UT, shows that the apparent rotation was due to the unfortunate juxtaposition of a star with the jet in rather poor seeing. There is no detectable rotation of the jet and this may, possibly, have been a renewed outbreak of the previous active region. On Sept. 26, we detected the p.a. of the end of the ‘spiral arm’ and not the ‘straight jet’ section. The expansion of the jet (which faded rapidly) and its evolution look very much like the earlier outburst. Projected distances of the linear section of the jet (up to the sharp kink) from our data: Sept. 28.86, 2′′.7 (12 800 km); Oct. 2.86, 4′′.9 (23 300 km). This implies an expansion velocity of 30 m/s, with jet commencement on Sept. 24.2. Images taken on Oct. 14 and 15 show yet another jet emerging, close to the nucleus.”

1995 October 16 Daniel W. E. Green